### v1

@@ 2.2 $p$ values
If we have a confidence interval, we may use it to define a $p$ value as follows: Let $C_{\alpha}(N)$ be a confidence set for $\theta$ with confidence level $1-\alpha$. Let, for any parameter value $\theta_{0}$,
$A_{\alpha}(\theta_{0})\equiv\{n|\theta_{0}\in C_{\alpha}(n)\}.$ (16)
@@
For a given observation $n$, let $p$ be the largest value of $\alpha$ such that $n\in A_{\alpha}(\theta_{0})$. That is, $A_{\alpha}(\theta_{0})$ is interpreted as the acceptance region, at the $\alpha$ significance, for the test $H_{0}:\theta=\theta_{0}$. The observed value of $n$ is just barely consistent with $H_{0}$ at significance level $p$ [14].
@@
A small value of $p$ corresponds to a large value of $1-\alpha$, i.e., to a confidence interval with a large confidence level. That is, we must have a large confidence interval in order to contain the observation $n$. The smaller $p$ is, the less probable our observation is under $H_{0}$. We would like to design our confidence interval such that we obtain $p$ values that can be interpreted as something like the probability of obtaining an observation as far or farther from the value $\theta_{0}+b$ as the observed value $n$. If we have a confidence interval, $C_{\alpha}(N)=(\ell_{\alpha}(N),u_{\alpha}(N))$, we may compute a table of $p$ values as follows: For the hypothesis $\theta_{0}=\ell_{\alpha}(N)$, $P(N|\theta_{0})=\alpha$ is the $p$ value for this test. And for the hypothesis $\theta_{0}=u_{\alpha}(N)$, $P(N|\theta_{0})=\alpha$ is the $p$ value for this test.
@@
A common case is for a null hypothesis that there is no signal $\hbox{$H_{0}$:\ }\theta_{0}=0$, and the alternative hypothesis of interest is that there is a signal, $\hbox{$H_{1}$:\ }\theta>0$. If we have devised a test for this, with acceptance region $A_{\alpha}(0)$, then again $p$ is computed as the largest value of $\alpha$ for which $n\in A_{\alpha}(0)$. Since this is a one-sided test, there is a unique interpretation that we would like to have: The most sensible $p$ value is one that gives the probability that $N\geq n$ under the hypothesis $\hbox{$H_{0}$:\ }\theta_{0}=0$, where $n$ is the observed value. That is,
$p=P(N\geq n|\theta=0).$ (17)
@@ 3 Desirable properties of confidence intervals
Over the decades, an extensive discussion on confidence intervals for discrete distributions including Poisson has developed in the statistics literature, with additional contributions from the physics community. The discreteness of the sample space has complicated the discussion, although many of the same issues hold as well in the case of continuous distributions. Various properties are considered desirable by one or another investigator. The situation in physics remains confused, with proponents of different algorithms and no consensus. In this paper we discuss what properties could be considered most important and then investigate several approaches for their properties and formulate recommendations.
There are no unknown nuisance parameters in this problem, so obtaining solutions to Eq. 6 is relatively straightforward. In spite of this, there are many approaches that have been suggested in the statistics literature, plus some in the physics community. Comparative reviews of several methods in the statistics literature appear, for example, in refs. [15], [16], [17], and [18]. Here, we discuss matters in the context of physics measurements. In doing so, we aren’t suggesting that physics measurement differs fundamentally from other fields, but we are interested in the context of the culture of thought in physical measurement.
@@ 3.2 Connectedness
There is no requirement in Eq. 6 that a confidence set be connected. For some distributions, disconnected sets provide useful descriptions. However, the Poisson distribution is unimodal, and our sets will be easiest to understand if we insist on connected confidence sets (e.g., [22]). That is, we will insist that the sets $C_{\alpha}(N)$ contain only single continuous ranges of $\theta$ values. This is not always automatic in practice, we must enforce it. This requirement will put constraints on our ability to minimize overcovering. With this requirement, our confidence set can be described as an interval:
$C_{\alpha}(N)=\left[\ell(N),u(N)\right].$ (18)
We henceforth require connected intervals. If we have a confidence set that is not connected, we can always add intermediate parameter values until it is connected. This will be at the cost of performance in terms of overcoverage.
@@ 3.3 Contains maximum likelihood estimator
For any observation, $n$, we require that the confidence interval $(\ell,u)$ for $\mu$ contains $n$: $\ell\leq n\leq u$. That is, the MLE for $\mu$ is in the confidence interval, and the confidence interval for $\theta$ is $\ell-b,u-b]$ is $\ell-b\leq n-b\leq u-b$, containing the MLE nbn-b for $\theta$. [^4] For technical reasons, occurring when a lower endpoint for some observation coincides with an upper endpoint for another observation, Casella and Robert [21] argue for half-open confidence intervals, $\ell(n),u(n))$. In this case our requirement for a $\theta$ confidence interval $\ell-b,u-b)$ is $\ell-b\leq n-b<u-b$. As in [21], we suspect violation of this will be non-optimal in other respects, and in any event, it is a descriptive property that we would normally expect and hence should probably require.
@@ 3.4 Optimal coverage
$\inf_{\theta}P\left\{\theta\in[\ell(N),u(N))|\theta\right\}=1-\alpha.$ (19)
That is, for at least some $\theta$, the confidence interval gives exact coverage. This is a reasonable property to strive for, since if we have an interval that doesn’t achieve this, it seems plausible that we could improve it until it does.
@@ 3.5 Length
$P(\theta\in C^{\prime}|\theta)$ $\geq P(\theta\in C|\theta),\quad\hbox{with inequality for some $\theta$.}$ (23)
@@
Short lengths are considered desirable to localize the description of the measurement. However, we will see that it must be done with some care that some values of $n$ don’t lead to absurdly short intervals. That is, we have intuition that the variance of the Poisson is $\mu$, and hence intervals should scale like $\sqrt{n}$ for different $n$ values. If we violate this, then our intervals lack the intuitive connection of a good description.
@@ 3.6 Scale and asymptopia
It should be easy to make an intuitive connection with the result. For Poisson sampling (Eq. 5), the variance of the distribution is $\theta+b$ and hence $\sqrt{\theta+b}$, or estimator $\sqrt{N}$, sets the scale we expect for the size of a confidence set. Indeed, an acceptable confidence set should most naturally asymptotically approach $(\hat{\theta}-f_{\alpha}\sqrt{\hat{\theta}+b},\hat{\theta}-f_{\alpha}\sqrt{\hat{\theta}+b})$, where $\hat{\theta}=N-b$ and $f_{\alpha}=\hbox{erf}^{-1}(1-\alpha)$.
@@ 3.8 Symmetry
Another property that might be desirable for a good description is a notion of symmetry. Here, following our frequency interpretation, symmetry is taken to mean that the interval is constructed with approximately equal probability content in the high and low tails outside the interval. We’ll call this “symmetric in probability”. This is indeed the classic Poisson interval, and we consider it in Sec. 5.1.
A different notion of symmetry is that of equal size intervals on both sides of the chosen point estimator. We’ll call this “symmetry in interval” to distinguish from the above notion of symmetry. In this case, the result of the measurement can be described as the point estimator “plus or minus cc”, where $c$ is one-half the confidence interval size (with confidence level traditionally chosen as 68% [^5] Motivated by the normal probability to be within one standard deviation of the mean. When we say 68% here, we always mean more precisely this normal probability.). While perhaps a nice feature, we do not have a good reason to insist on it, although we’ll use such intervals as convenient in our discussion of averages in Section 9.
We aren’t interested in searching explicitly for either one-sided or two-sided intervals, leaving this property to come naturally out of the algorithm. For the descriptive nature of frequentist statistics, it is more important to achieve a good description and insisting on two-sided or one-sided intervals is not helpful. Averaging of results from different measurements can be awkward with one-sided intervals. We discuss limits further in Section 4. On the other hand, one-sided intervals, especially upper limits, are often very useful in interpretive statements, such as “I am 95% sure that the true value is less than five.” This is a degree-of-belief statement, forming the basis for making a decision, and is the domain of Bayesian analysis. While we will investigate the frequency properties of two choices of uninformative prior in Section 7, Bayesian analysis is outside the scope of our present interest.

### v2

@@ 2.2 $p$-values
If we have a confidence interval derived from a hypothesis test, we may use the acceptance regions $A_{\alpha}(\theta_{0})$ to obtain a $p$-value as follows: Let, for any parameter value $\theta_{0}$,
$A_{\alpha}(\theta_{0})\equiv\{n|\theta_{0}\in C_{\alpha}(n)\},$ (16)
@@
be the acceptance set corresponding to confidence interval $C_{\alpha}(n)$. For a given observation $n_{\rm obs}$, if $n_{\rm obs}<\theta_{0}$, find the smallest $\alpha$ such that the smallest element of $A_{\alpha}(\theta_{0})$ is $n_{\rm obs}$. If $n_{\rm obs}>\theta_{0}$, find the smallest $\alpha$ such that the largest element of $A_{\alpha}(\theta_{0})$ is $n_{\rm obs}$. Then $p=\alpha$. That is, $A_{\alpha}(\theta_{0})$ is interpreted as the acceptance region, at the $\alpha$ significance, for the test $H_{0}:\theta=\theta_{0}$. The observed value of $n$ is just barely consistent with $H_{0}$ at significance level $p$ [38].
@@
A small value of $p$ corresponds to a large value of $1-\alpha$, i.e., to a confidence interval with a large confidence level. That is, we must have a large confidence interval in order to contain the observation $n$. The smaller $p$ is, the less probable our observation is under $H_{0}$. We would like to design our confidence interval such that we obtain $p$-values that can be interpreted as something like the probability of obtaining an observation as far or farther from the value $\theta_{0}+b$ as the observed value $n$.
@@
A common case is for a null hypothesis that there is no signal $\hbox{$H_{0}$:\ }\theta=\theta_{0}=0$, and the alternative hypothesis of interest is that there is a signal, $\hbox{$H_{1}$:\ }\theta>0$. If we have devised a test for this, with acceptance region $A_{\alpha}(0)$, then again $p$ is computed as the smallest value of $\alpha$ for which $n\in A_{\alpha}(0)$. Since this is a one-sided test, there is a unique interpretation that we would like to have: The most sensible $p$-value is one that gives the probability that $N\geq n$ under the hypothesis $\hbox{$H_{0}$:\ }\theta_{0}=0$, where $n$ is the observed value. That is,
$p=\hbox{Prob}(N\geq n|\theta=0).$ (17)
3 Descriptive statistics
The first draft of this paper received a comment from Cousins [8] on two main issues:
  1. 1.
    I have suggested above that the purpose of frequency statistics is descriptive, that is, it makes statements about the result of the observed sampling. Cousins suggests that it also satisfies an inferential purpose.
  2. 2.
    I have suggested in the discussion below Eq. 2 that, when describing the result of a sampling, the likelihood function may sometimes be usefully evaluated at parameter values that are known to be incorrect (e.g., “non-physical”). Cousins suggests that this is forbidden.
I consider these two issues to be related. I have posted a response [31] but it seems desirable to include an expanded discussion in published form here.
We first address the descriptive vs. inferential property. Kendall, Stuart, Ord, and Arnold define “statistical inference” as an “inductive process from sample to population” (section 26.1 of [43]). Shao defines “statistical inference” in a somewhat self-referential way as making “an inference about the unknown population based on the sample $X$ (section 2.4 of [38]). Thus, the notion of inference involves making some statement about the sampling distribution based on an observed sample. Our Poisson case may be regarded as an example. In this case, the specific quantity of the population that we would like to infer is the population parameter $\theta$, or equivalently $\mu=\theta+b$. We note the we are not considering finite populations that might be of interest in other fields. That is, our inference problem is to make some statement about the Poisson mean given a sampling from the distribution.
The problem of inference (and induction) has a long history of philosophical thought and controversy. Karl Popper, for example, grapples with the problem of going from observations to statements about truth and falsity [30]. He considers the problems of inductive logic to be “insurmountable”, and that the problems “inherent in the doctrine …that inductive inference, although not ‘strictly valid’, can attain some degree of ‘reliability’ or of ‘probability’” (emphasis his) are likewise insurmountable. [^4] A recent essay on these matters appears in [16]. I subscribe to this point of view. That is, in my interpretation, statements about truth always require some level of subjectivity, which can be encoded as “degree of belief”. In other words, our knowledge results from a process of abductive reasoning.
Statistics philosopher Deborah Mayo and statistician David Cox published an article entitled “Frequentist statistics as a theory of inductive inference” [23]. In spite of the suggestive title, they do not argue that frequentist statistics provides an inference about truth, such as the value of a population parameter. Instead, frequentist statistics provide “evidence” that may be useful in forming an inference. This is precisely the distinction that I am making. A very small $p$-value under some null hypothesis ($H_{0}$) presents evidence against the null hypothesis, but must be taken in the context of all other available evidence before making the inference that $H_{0}$ is most likely incorrect. For example, if there is no other relevant evidence or experience, a $p$-value of 0.001 might be construed as strong evidence against $H_{0}$, but if this same $p$-value is the last in a sequence of 1000 results with an approximately uniform distribution of $p$-values, then the evidence against $H_{0}$ is weak.
The classic “Advanced Theory of Statistics” text [43] uses the expression “frequentist approach (e.g., section 26.4) in their chapter (26) entitled “Comparative Statistical Inference”. As far as I know, this reference avoids the expression “frequentist inference”, which I consider to be an astute choice, as I argue that frequentist statistics may be employed in forming an inference, but does not itself provide the inference. Their chapter 26 considers several approaches to estimation, including Bayesian. They recognize the controversial nature of the subject, and present their personal views in sections 26.58 through 26.78. I consider that my point of view is consistent with their discussion. In particular, their section 26.62 states:
“The frequency approach leads then to point and interval estimates and tests of hypotheses that are keyed to an interpretation of performance in the long run [italics theirs]. The other approaches we have described constitute several attempts to develop an additional, or alternative, notion of probability that enables the investigator to make inferential statements conditionally upon the data recorded in a particular statistical experiment [italics theirs].”
In the context of statistical methods for particle physics, Solmitz [^5] Solmitz is credited with introducing the maximum likelihood method to particle physics [33]. [39] describes both Bayesian and frequentist (“anti-Bayesian”) estimation, making the point in the latter methodology that no statement is actually made concerning the true value of a parameter without making a further step with Bayes’ principle. James and Roos [18] discuss reporting of results near an unphysical region, again noting that the frequentist statistics make no statement about the true value of a parameter.
People have sometimes proven that frequency statistics is “wrong”, resulting when they treat it as serving an inferential function. For one fairly recent prominent example, Trafimow and Marks [44], as editors of the journal Basic and Applied Social Psychology (BASP) famously banned the use of frequentist statistics such as $p$-values and confidence intervals from publication in their journal. They claimed that these statistics are “invalid”. The essence of the argument is contained in the abstract of [45]:
[…]
Here, by “standard null-hypothesis significance testing procedure” they mean frequency statistics such as $p$-values. The first sentence is a well-known observation. The second sentence is an assertion that frequency statistics provides an inference, which is incorrect.
It is remarkable that the editors state that no inferential procedures are required in BASP, but that “BASP will require strong descriptive statistics…”. By inferential procedures, they explicitly mean frequentist statistics with a milder stance on Bayesian statistics. Instead of embracing the descriptive frequentist statistics they go against their own requirement and ban them by misunderstanding them as inferential. In their defense, their ban is the result of finding that too many practitioners appear to also misunderstand the distinction and actually use $p$-values towards making stronger inferential conclusions than justified. This is the argument that people who do understand the distinction sometimes make that frequency statistics are problematic because they are often misunderstood (see, e.g., [2]). I prefer to think that education can mitigate this difficulty, and that we can benefit from use of frequency statistics.
The BASP editorial banning frequency statistics generated an immediate reaction (and some derision), with some people agreeing and others pointing out the flawed understanding. The American Statistical Association felt compelled to issue a statement clarifying what $p$-values are and aren’t and how they should be used [50]. This statement contains the important point about $p$-values: “It is a statement about data in relation to a specified hypothetical explanation, and is not a statement about the explanation itself.” That is, it is a description of the data in a relevant context, not an inference about the context.
I have explained the rationale for treating frequency statistics as descriptive rather than inferential. As noted, this does not mean that frequency statistics cannot be used in an inferential process. They just cannot be treated as themselves providing an inference. Instead, as a description of the results of a measurement, they provide evidence upon which conclusions about the parent population may be reached. I suggest that such conclusions in physics can best be expressed in a degree-of-belief paradigm.
Let us turn to the second point of contention, the suggestion that it is forbidden to evaluate the likelihood function in an unphysical region. It is useful here to make the distinction between a descriptive methodology and a degree-of-belief methodology. We will mostly use the example of sampling from the Poisson distribution with mean $\mu=\theta+b$ where $b$ is a known background parameter and $\theta$ is an unknown signal parameter, although the discussion is by no means limited to this example. We suppose that on physical grounds we know that $\theta\geq 0$. It is important to keep in mind that the sampling distribution and the likelihood function are not the same thing – the former gives probabilities for values of a random variable, given any parameter value(s), and the latter provides the value of the sampling distribution for a given value of the random variable, as a function of parameter value(s).
We may start with degree-of-belief, by which I mean that we wish to make statements about the true value of $\theta$ reflecting our belief (which may be different for different people). As we have complete belief that $\theta\geq 0$, we will never quote a result that includes negative values of $\theta$. In this case, we might as well say that the likelihood function is zero, or not defined, for $\theta<0$. It does not really matter, because we will always impose the $\theta\geq 0$ restriction on our statements of belief. In the context of using Bayes theorem, our prior belief is $P(\theta<0)=0$ and that imposes this constraint independent of what we do with the likelihood function.
The situation is quite different if we are describing our observation (measurement, or sampling from the parent distribution). In this case there is no obvious restriction because we are not trying to make a statement about the value of $\theta$. We are interested in providing a useful description of the result of the measurement.
It may be debated what constitutes a useful description, and I list several possibly desirable properties, not all mutually compatible, in section 4. But at the very least, it seems clear that we should provide a description that gives as much information contained in the measurement as possible, relevant to the question we are trying to address. In our Poisson example, this question is the value of $\theta$. That is, a description should include a sufficient statistic, if doing so is realistic. In our simple Poisson case, this is readily accomplished. Simply quoting the observed counts, $n$, is such a statistic. It is hard to imagine anyone objecting to quoting this statistic as the result of the sampling. Yet it is precisely the maximum likelihood estimator for $\mu$, if the likelihood is allowed to be evaluated in the non-physical region, $\mu<b$. If this is not to be permitted, the statistic becomes $\max(n,b)$. The corresponding point estimates for $\theta$ are nbn-b or $\max(n-b,0)$. If the likelihood function cannot be evaluated for $\theta<0$ (or $\mu<b$), the maximum likelihood estimator loses the sufficiency property, i.e., suppresses information. Allowing the likelihood function to be evaluated at negative $\theta$ maintains the sufficiency property of the point estimator, which I consider to be a strong argument for doing so.
Others also suggest that the likelihood function can be evaluated without constraining to physical parameter space. Solmitz [39] allows the likelihood function to be defined outside of the physical region, referring to the position of the maximum as “which may be outside of the physical region”. James and Roos [18] argue that when publishing the description of the result of a measurement “it should even be as much as possible independent of physical constraints. The main reason is that only then can the results be unbiased and therefore easily combined with other results.”
Unbiasedness is generally regarded as a desirable property in a point estimator, and nbn-b is an unbiased estimator for $\theta$. The property of smallest error, as in the mean squared error (MSE) is also a desirable property and sometimes we wish to use such a statistic (especially in trading off bias and variance in the context of density estimation). For this property, the biased estimator $\max(n-b,0)$ has a smaller MSE than nbn-b. However, if we have nbn-b we can also form $\max(n-b,0)$; we cannot go in the other direction, as we have lost sufficiency.
A familiar situation arises when measuring a signal above a (here assumed to be known) background distribution in some sampled distribution expressed as a histogram with $n$ bins with contents $x=\{x_{1},\ldots,x_{n}\}$. If the statistics of the background is large, the Poisson bin contents are approximately normally distributed. Assume for simplicity here that the shape and location of the signal pdf is known and the only unknown is the signal strength. Suppose that we do a maximum likelihood fit to estimate the signal strength. We wish, for example, to test the goodness of fit of our model to the data. Let $L_{\rm max}$ be the maximum likelihood for our fit to the data according to the model with signal strength varied. Let $L_{0}$ be given by
$L_{0}=\prod_{i=1}^{n}\frac{{x_{i}}^{x_{i}}}{x_{i}!}e^{-x_{i}}.$ (18)
Let $\lambda=-2\ln L_{\rm max}/L_{0}$. In the asymptotic region, the goodness of fit of the data to the signal plus background model is obtained according to $\lambda$ being approximately $\chi^{2}$ distributed with n1n-1 degree of freedom. This statement holds as long as the signal strength is allowed to be negative as well as positive, i.e., $L_{\max}$ is to be computed without constraining the signal to be in the physical region (to satisfy the interior of the maintained hypothesis according to Wilk’s theorem). While we can investigate the distribution of $\lambda$ with a non-negative signal constraint, it is certainly convenient to be able to use the $\chi^{2}$ distribution, as is commonly done. This provides another situation supporting the evaluation of likelihoods in the non-physical region.
In the context of descriptive statistics, we have seen several ways in which it is useful to evaluate the likelihood function in non-physical regions. We know of no reason why this should be forbidden in this context. There is no reason to evaluate the likelihood function in a non-physical rtegion in the context of degree-of-belief statistics however, and perhaps this is the source of some resistance to the idea.
@@ 4 Desirable properties of confidence intervals
Over the decades, an extensive discussion on confidence intervals for discrete distributions including Poisson has developed in the statistics literature, with additional contributions from the physics community. The discreteness of the sample space has complicated the discussion, although many of the same issues hold as well in the case of continuous distributions. Various properties are considered desirable by one or another investigator. Unfortunately, these “desirable” properties cannot all be optimally satisfied at once and compromises must be made. The situation in physics remains confused, with proponents of different algorithms and no consensus. In this paper we discuss what properties could be considered desirable and then investigate the properties of several approaches and formulate recommendations.
There are no unknown nuisance parameters in this problem, and obtaining solutions to Eq. 6 is relatively straightforward without having to deal with nuisance parameters. Indeed, there are many approaches that have been suggested in the statistics literature, plus some in the physics community. Comparative reviews of several methods in the statistics literature appear, for example, in refs. [5], [17], [28], and [1]. Here, we discuss matters in the context of physics measurements. In doing so, we aren’t suggesting that physics measurement differs fundamentally from other fields, but we are interested in the context of the culture of thought in physical measurement.
@@ 4.2 Connectedness
There is no requirement in Eq. 6 that a confidence set be connected. For some distributions, disconnected sets provide useful descriptions. However, the Poisson distribution is unimodal, and our sets will be easiest to understand if we insist on connected confidence sets (e.g., [4]). That is, we will insist that the sets $C_{\alpha}(N)$ contain only single continuous ranges of $\theta$ values. This is not always automatic in practice; we must enforce it. This requirement will put constraints on our ability to minimize overcovering. With this requirement, our confidence set can be described as an interval:
$C_{\alpha}(N)=\left[\ell_{\alpha}(N),u_{\alpha}(N)\right].$ (19)
We henceforth require connected intervals. If we have a confidence set that is not connected, we can always add intermediate parameter values until it is connected. This will be at the cost of performance in terms of overcoverage. For simplicity we will also drop the $\alpha$ subscript on $\ell$ and $u$ from now on as understood.
@@ 4.3 Contains maximum likelihood estimator
For any observation, $n$, we require that the confidence interval $(\ell,u)$ for $\mu$ contains $n$: $\ell\leq n\leq u$. That is, the MLE for $\mu$ is in the confidence interval, and the confidence interval for $\theta$ is $\ell-b,u-b]$ is $\ell-b\leq n-b\leq u-b$, containing the MLE nbn-b for $\theta$. [^6] For technical reasons, occurring when a lower endpoint for some observation coincides with an upper endpoint for another observation, Casella and Robert [6] argue for half-open confidence intervals, $\ell(n),u(n))$. In this case our requirement for a $\theta$ confidence interval $\ell-b,u-b)$ is $\ell-b\leq n-b<u-b$. As in [6], we suspect violation of this will be non-optimal in other respects, and in any event, it is a descriptive property that we would normally expect (it seems bizarre to quote a confidence interval that excludes the point estimate) and hence should probably require. As discussed in section 3, we insist that the maximum likelihood estimator not be restricted to some physical parameter space.
@@ 4.4 Optimal coverage
$\inf_{\theta}\hbox{Prob}\left\{\theta\in[\ell(N),u(N)]|\theta\right\}=1-\alpha.$ (20)
That is, for at least some $\theta$, the confidence interval gives exact coverage. This is a reasonable property to strive for, since if we have an interval that does not achieve this, it seems plausible that we could improve it until it does.
@@ 4.5 Length
$\hbox{Prob}(\theta\in C^{\prime}|\theta)$ $\geq\hbox{Prob}(\theta\in C|\theta),\quad\hbox{with inequality for some $\theta$.}$ (24)
@@
Short lengths are considered desirable to localize the description of the measurement. However, we will see that it must be done with some care that some values of $n$ do not lead to absurdly short intervals. That is, we have intuition that the variance of the Poisson is $\mu$, and hence intervals should scale like $\sqrt{n}$ for different $n$ values. If we violate this, then our intervals lack the intuitive connection of a good description.
@@ 4.6 Scale and asymptopia
It should be easy to make an intuitive connection with the result. For Poisson sampling (Eq. 1), the variance of the distribution is $\theta+b$ and hence $\sqrt{\theta+b}$, or estimator $\sqrt{N}$, sets the scale we expect for the size of a confidence set. Indeed, an acceptable confidence set should most naturally asymptotically approach $(\hat{\theta}-f_{\alpha}\sqrt{\hat{\theta}+b},\hat{\theta}-f_{\alpha}\sqrt{\hat{\theta}+b})$, where $\hat{\theta}=N-b$ and $f_{\alpha}=\sqrt{2}\,\hbox{erf}^{-1}(1-\alpha)$.
@@ 4.8 Symmetry
Another property that might be desirable for a good description is a notion of symmetry. Here, following our frequency interpretation, symmetry is taken to mean that the interval is constructed with approximately equal probability content in the high and low tails outside the interval. We will call this “symmetric in probability”. This is indeed the classic Poisson interval, and we consider it in section 6.1.
A different notion of symmetry is that of equal size intervals on both sides of the chosen point estimator. We will call this “symmetry in interval” to distinguish from the above notion of symmetry. In this case, the result of the measurement can be described as the point estimator “plus or minus cc”, where $c$ is one-half the confidence interval size (with confidence level traditionally chosen as 68% [^7] Motivated by the normal probability to be within one standard deviation of the mean. When we say 68% here, we always mean more precisely this normal probability.). While perhaps a nice feature, we do not have a good reason to insist on it, although we will use such intervals as convenient in our discussion of averages in section 10.
We aren’t interested in searching explicitly for either one-sided or two-sided intervals, leaving this property to come naturally out of the algorithm. For the descriptive nature of frequentist statistics, it is more important to achieve a good description and insisting on two-sided or one-sided intervals is not helpful. Averaging results from different measurements can be awkward with one-sided intervals. We discuss limits further in section 5. On the other hand, one-sided intervals, especially upper limits, are often very useful in interpretive statements, such as “I am 95% sure that the true value is less than five.” This is a degree-of-belief statement, forming the basis for making a decision, and is the domain of Bayesian analysis. While we will investigate the frequency properties of two choices of uninformative prior in section 8, Bayesian analysis is outside the scope of our present interest.